# Droplets sliding on soft solids shed elastocapillary rails

### v1

@@ I Introduction
Figure 1: The platform for studying droplet motion on soft substrates. (a) Side view of the droplet between the capillary and the substrate; see also in Movie 1. (b) Schematic of the experimental setup. (c) The measured force $F$ as a function of the droplet velocity $v$. The colored shadow represents the standard deviation across three experiments on different samples.
@@ II.2 Force and Deformation Measurements
We employ a Linnik interference imaging setup to quantify soft substrate deformation [25, 21]. An LED light source ($\lambda=643$ nm) illuminates the substrate from below, and reflected light from the water-substrate interface interferes with a reference beam. Figure 2 displays typical captured patterns. These interference patterns, when processed with phase retrieval methods, enable high-resolution analysis of substrate deformation on the order of 1 nm (details and demo code in ESI). Additionally, the interferometric imaging simultaneously captures droplet shapes (dark shadows, Fig. 2) and capillary positions (shallow shadows, Fig. 2).
@@ III.1 Force measurements
We compare these results to a viscoelastic model of an isotropically spreading droplet on a deformable substrate [10]. After substituting the wetting and rheological properties of our softer gel into the model, and twice the resting diameter of the droplet as an approximation for the length of the contact line (details in ESI; ‘twice’ is to account for both advancing and receding fronts, with further discussion of the choice of the contact line size in Section III.2), we obtain a theoretical prediction for the drag on the droplet on the softer gel, shown as dashed red line in Fig. 1(c). This prediction overestimates the experimentally measured drag.
@@ III.2 Droplet shapes
While the model prediction in Section III.1 conveniently used twice the resting droplet diameter as an approximation for contact line size, Refs. 18, 20 emphasize that droplets can undergo significant shape changes during sliding. Our experiments confirm this, especially at higher velocities on softer substrates. Typical droplet shapes captured by interferometric imaging during sliding from left to right are shown in Fig. 2 (dark shadows represent the droplets).
At lower velocities, droplets on both softer and stiffer gels maintain a nearly circular contact line [Figs. 2(a, c)]. However, at higher velocities, the contact line elongates in the direction of motion and narrows in the transverse direction [Figs. 2(b, d)]. This reshaping is particularly pronounced on softer gels, where the droplets take on a bullet-like form: The contact lines become straight and parallel to the direction of motion. We quantify this reshaping by measuring the droplet’s length, $l$, and width, $w$ [Fig. 2(e)]. The aspect ratio, $l/w$, increases with velocity on both gels, with a larger shape change on the softer gel. This increase in the aspect ratio is mainly driven by the change in the droplet width, as the change in the length of the droplet is minor (detailed measurements in Fig. S4 in ESI).
@@
Figure 3: The measured force per unit width of the moving droplet as a function of velocity $v$. The colored shadow represents the standard deviation across three experiments.
@@
Figure 4: The relaxation of the elastocapillary rails. (a-c) Interferometric images showing the time evolution of the deformed substrate: $t=$ 0 s (a), 2 s (b), and 10 s (c). The corresponding movie is Movie 6. (d) The relaxation of the substrate ($h$) along the vertical dashed line in (a-c). (e) Zoom-in profile of the rails near their tips. The rail shape is normalized by its peak height, $h_{\mathrm{peak}}$. The inset shows the rail profile (height difference from the peak, $h-h_{\mathrm{peak}}$) as a function of $y-y_{\mathrm{peak}}$. (f) Relaxation of the rail height $h_{\mathrm{peak}}$ as a function of the length along the rail, $x$. (g) Relaxation of the rail height $h_{\mathrm{peak}}$ over time, $t$. The colored shadow shows the standard deviation (details in Fig. S5 in ESI).
@@
Despite the confinement in the narrow gap between the capillary and the substrate, we observe a decrease in droplet width with increasing velocity (detailed measurements in Fig. S5 in ESI). This suggests that using dynamic droplet width, rather than its size at rest, is a more accurate way to characterize the size of the contact line involved in energy dissipation. Accordingly, we calculate the drag per unit width by dividing the measured force by twice the droplet width, Fig. 3. By considering the droplet shape change, the experimental data moves closer (but certainly not collapsing) to the 2D linear viscoelastic model prediction (dashed red line in Fig. 3), particularly at low velocities.
A key difference between our experiment and the 2D model is that the model assumes straight contact lines perpendicular to the direction of motion. Sliding droplets, however, involve curved contact lines with a significant velocity component parallel to them. Nevertheless, the existing 2D model provides satisfactory order-of-magnitude estimations for dissipation [22]. Simple extensions of this model accounting for finite substrate thickness and circular contact line shape (ESI) result in a negligible change to the predicted dissipation.
@@ III.3 Elastocapillary Rails
At high speeds, droplets not only assume a bullet-like morphology, but they also leave behind an unusual pair of parallel rails. These structures are clearly visible in our interferometric images [Figs. 2(c, e), Fig. 4(a), and Movies 5 and 6)]. They appear as a pair of parallel deformations to interferometric fringes following the direction of the droplet motion, and connecting to its sides. These fringes indicate surface topography, which we quantify using phase retrieval methods described in Refs. 26, 27, 28 (details in ESI). This method achieves a remarkably high out-of-plane resolution on the order of 1 nm. The resulting profile, shown in Fig. 4(d), reveals a pair of parallel ridges we term ‘elastocapillary rails.’ These rails can reach up to $O(1)$ mm in length. They are tallest near the droplet, with a height of $O(100)$ nm from the rear of the droplet. Interferometry does not allow measurements closer to the contact line because steep profiles prevent resolution of the fringes.
The shape of the elastocapillary rails shares some features with typical wetting ridges, in the absence of a contact line. Analyzing the deformed surface shape along a vertical line in the interferometric images [Figs. 4(a-c)] reveals their cross-sectional profile [Fig. 4(d)]. They have a width of about $50$ $\mu$m, compared to the elastocapillary length of $\gamma/E\approx 10$ $\mu$m, and ‘dimples’ on either side that are known to appear when the substrate thickness becomes comparable to the elastocapillary length [29]. In contrast to wetting ridges, however, the tips of the rails are rounded [30, 21] rather than triangular [29]. Notably, the rails extend beyond the immediate vicinity of the contact line and detach from the droplets. Finally, the shape of the tips of these wetting ridges evolves over time in a self-similar fashion. If we shift the tips of the rails at different times, we see that their shapes do not overlap [inset in Fig. 4(e)]. Instead, the slopes along the sides and the curvatures at the tip decrease over time or distance from the droplet. However, scaling the rail cross-section by its height collapses the profiles at different locations/times [Fig. 4(e)]. Self-similar evolution is consistent with the viscoelastic relaxation model [10].
@@ IV Discussion and Conclusions
Our interferometric platform enabled simultaneous investigation of multiscale structure and dynamics of droplets sliding on a deformable substrate. At higher droplet velocities, droplets adopt a distinct bullet-like shape, the drag force levels off, and prominent parallel elastocapillary rails are shed from the droplet. Feedback between the microscopic structure of the elastocapillary ridge and the overall shape of the droplet plays an essential role in these phenomena. Sliding along a wetting ridge creates much less drag than surfing across it. This favors the narrowing of the droplet, creating longer droplets. At high speeds, the length of the parallel side contact lines of the bullet-like droplets is similar to the droplet length, $l$.
Points along the sides of a sliding droplet are pulled upward by the liquid-vapor interface for a time $O(l/v)$ while points at the front and back of the droplet are only pulled upward for a time $O(\gamma/Ev)$. Thus, the ridges on the side of the droplet are loaded by a factor of $EL/\gamma=O(100)$ longer than the front. Previous work [31, 21] shows that poroelasticity can lead to a ridge decay-time comparable to the time of contact between the droplet and substrate. Thus, elastocapillary ridges formed along the direction of motion would be more persistent, than those at the leading and trailing edges of the droplet. This leads to the formation of sheddable elastocapillary rails. The relaxation data in Fig. 4(g) supports this hypothesis, as rails formed a low speeds (i.e., long contact times) take longer to decay.

### v2

@@ I Introduction
Figure 1: The platform for studying droplet motion on soft substrates. (a) Side view of the droplet between the capillary and the substrate; see also in Movie 1. (b) Schematic of the experimental setup. (c) The measured force $F$ as a function of the droplet velocity $v$. The dashed line represents the prediction of a 2D, linear, viscoelastic model for a moving, circular contact line, with a diameter equal to that of the resting droplet. The colored shadow represents the standard deviation across three experiments on different samples.
@@ II.2 Force and Deformation Measurements
We employ a Linnik interference imaging setup to quantify soft substrate deformation [25, 21]. An LED light source ($\lambda=643$ nm) illuminates the substrate from below, and reflected light from the water-substrate interface interferes with a reference beam. Figure 2 displays typical captured patterns. These interference patterns, when processed with phase retrieval methods, enable high-resolution analysis of substrate deformation on the order of 1 nm (details and demo code in ESI). Additionally, the interferometric imaging simultaneously captures droplet shapes (dark shadows, Fig. 2) and capillary positions (light shadows, Fig. 2).
@@ III.1 Force measurements
This dissipation can arise either from the solid substrate, the contact line, or the liquid droplet. The capillary number, $\mathrm{Ca}=\mu v/\gamma$, where the viscosity $\mu\approx 1\times 10^{-3}$ Pa s, characterizes the relative importance of viscosity and surface tension in contact-line motion [26, 27]. In our experiments, Ca varies from $10^{-8}$ to $10^{-5}$, indicating that viscous flow effects are likely negligible [28, 29, 30]. Furthermore, our measurements [Fig. 1(c)] show that the drag on the hard gel is much smaller than on the soft gel. This suggests that dissipation due to viscosity as well as contact-line pinning is a minor contributor to the total dissipation, particularly at low velocities.
Thus, the dissipation must occur in the substrate. We compare the measured drag force to a linear viscoelastic model, adapted from Ref. 10, which originally described an isotropically spreading droplet on a thin, deformable substrate. We approximate the experimental system as a moving circular contact line with a diameter matching that of the resting droplet. We estimate the drag force per length at each point on the contact line by decomposing the motion onto vectors perpendicular and parallel to the contact line. We assume the component of motion along the contact line generates no drag and the component perpendicular to the contact line generates the same resistance as an isotropically expanding droplet. The total drag force is then obtained by integrating over the entire circular contact line. Substituting the wetting and rheological properties of our softer gel into the model yields a theoretical prediction, shown as the dashed red line in Fig. 1(c). Notably, this prediction overestimates the experimentally measured drag. Further details of the model and calculations are available in the ESI. In the next Section, we will demonstrate that a part of this overestimation is related to the change in the droplet shape.
@@ III.2 Droplet shapes
While the simple model in Section III.1 conveniently assumed that the droplet shape was unchanged by motion, Refs. 18, 20 emphasize that droplets can undergo significant shape changes during sliding. Our experiments confirm this, especially at higher velocities on softer substrates. Typical droplet shapes captured by interferometric imaging during sliding from left to right are shown in Fig. 2 (dark shadows represent the droplets).
@@
Figure 3: The measured (a) width $w$ and (b) force per unit width of the moving droplet as a function of velocity $v$. The dashed line represents the prediction of a 2D, linear, viscoelastic model for a moving, circular contact line. The colored shadow represents the standard deviation across three experiments.
At lower velocities, droplets on both softer and stiffer gels maintain a nearly circular contact line [Figs. 2(a, c)]. However, at higher velocities, the contact line elongates in the direction of motion and narrows in the transverse direction [Figs. 2(b, d)]. This reshaping is particularly pronounced on softer gels, where the droplets take on a bullet-like form: The contact line on the side of the droplet straightens, parallel to the direction of motion. We quantify this shape-change by measuring the droplet’s length, $l$, and width, $w$ [Fig. 2(e)]. The aspect ratio, $l/w$, increases with velocity on both gels, with a larger shape change on the softer gel. This increase in the aspect ratio is mainly driven by the change in the droplet width, as the change in the length of the droplet is minor [measurements in Fig. 3(a) and Fig. S4 in ESI]. Note that the contact line is expected to remain circular in the limit of low speeds, as the net dissipative force vanishes [10]. However, this regime has not been reached in our experiments.
@@
Figure 4: The relaxation of the elastocapillary rails. The substrate is the softer gel ($E=4$ kPa), and the velocity of the sliding droplet is $v=1$ mm/s. (a-c) Interferometric images showing the time evolution of the deformed substrate: $t=$ 0 s (a), 2 s (b), and 10 s (c). The corresponding movie is Movie 6. (d) The relaxation of the substrate ($h$) along the vertical dashed line in (a-c). (e) Zoom-in profile of the rails near their tips. The rail shape is normalized by its peak height, $h_{\mathrm{peak}}$. The inset shows the rail profile (height difference from the peak, $h-h_{\mathrm{peak}}$) as a function of $y-y_{\mathrm{peak}}$. (f) Relaxation of the rail height $h_{\mathrm{peak}}$ as a function of the length along the rail, $x$. (g) Relaxation of the rail height $h_{\mathrm{peak}}$ over time, $t$. The colored shadow represents the standard deviation, calculated from 737 measurements taken at different positions along the rail within the interferometry image sequences [i.e., different x as shown in (b)] from a single experiment. Further details are provided in Fig. S6 in the ESI.
@@
Despite confinement in the narrow gap between the capillary and the substrate, we observe a decrease in droplet width with increasing velocity, shown in Fig. 3(a). This suggests that using dynamic droplet width, rather than its size at rest, is a more accurate way to characterize the size of the contact line involved in energy dissipation. Accordingly, we calculate the drag per unit width by dividing the measured force by twice the droplet width, Fig. 3(b). By considering the droplet shape change, the experimental data moves closer (but certainly not collapsing) to the 2D linear viscoelastic model prediction [dashed red line in Fig. 3(b)], particularly at low velocities.
The model still overestimates dissipation at high velocities, which is likely due to several simplifications. The model assumes a linear viscoelastic response, while nonlinear effects may become significant at large deformations [7]. Additionally, the model integrates contributions from independent points on the circular contact line, treating the contact line as locally straight and neglecting the potential influence of its overall curvature. We also note that we treat the surface tensions of the solid substrate and liquid as constants [21], while it is known that the surface tension of an aqueous liquid can decrease in contact with a silicone gel [23], and the surface tension of a soft solid can vary significantly under deformation [31, 5, 32].
@@ III.3 Elastocapillary Rails
At high speeds, droplets not only assume a bullet-like morphology as previously described, but they also leave behind a striking pair of parallel rails. These structures are clearly visible in our interferometric images on softer gels [Figs. 2(c, e), Fig. 4(a), and Movies 5 and 6)]. Such rails are not observed on the stiffer elastomer. They appear as a pair of parallel deformations to interferometric fringes following the direction of the droplet motion, and connecting to its sides. These fringes indicate surface topography, which we quantify using phase retrieval methods described in Refs. 33, 34, 35 (details in ESI). This method achieves a remarkably high out-of-plane resolution on the order of 1 nm. The resulting profile, shown in Fig. 4(d), reveals a pair of parallel ridges we term ‘elastocapillary rails.’ These rails can reach up to $O(1)$ mm in length. They are tallest near the droplet, with a height of $O(100)$ nm from the rear of the droplet. Interferometry does not allow measurements closer to the contact line because steep profiles prevent resolution of the fringes.
The shape of the elastocapillary rails shares some features with typical wetting ridges, in the absence of a contact line. Analyzing the deformed surface shape along a vertical line in the interferometric images [Figs. 4(a-c)] reveals their cross-sectional profile [Fig. 4(d)]. They have a width of about $50$ $\mu$m, compared to the elastocapillary length of $\gamma/E\approx 10$ $\mu$m, and ‘dimples’ on either side that are known to appear when the substrate thickness becomes comparable to the elastocapillary length [36]. In contrast to wetting ridges, however, the tips of the rails are rounded [37, 21] rather than triangular [36]. Notably, the rails extend beyond the immediate vicinity of the contact line and detach from the droplets. Finally, the shape of the tips of these wetting ridges evolves over time in a self-similar fashion. If we shift the tips of the rails at different times, we see that their shapes do not overlap [inset in Fig. 4(e)]. Instead, the slopes along the sides and the curvatures at the tip decrease over time or distance from the droplet. However, scaling the rail profile by its peak height collapses the profiles at different locations/times [Fig. 4(e)]. Self-similar evolution is consistent with the viscoelastic relaxation model [10].
@@ IV Discussion and Conclusions
Our interferometric platform enabled simultaneous investigation of multiscale structure and dynamics of droplets sliding on a deformable substrate. At higher droplet velocities, droplets adopt a distinct bullet-like shape, the drag force levels off, and prominent parallel elastocapillary rails are shed from the droplet.
In contrast to our previous work with similar-sized droplets sliding under gravity [15], where the bullet-like shape indicative of elastocapillary rails was not observed, the droplets in our current study move at speeds approximately 100 times faster. This suggests that elastocapillary rails for millimeter-sized droplets are likely limited to strongly driven systems, such as the shear geometry employed here or potentially in a spin coater.
Feedback between the microscopic structure of the elastocapillary ridge and the overall shape of the droplet plays an essential role in these phenomena. Sliding along a wetting ridge creates much less drag than surfing across it. This favors the narrowing of the droplet, creating longer droplets. At high speeds, the length of the parallel side contact lines of the bullet-like droplets is similar to the droplet length, $l$.
We hypothesize that elastocapillary rails are shed at high speeds because points along the sides of the droplet are pulled upward for about 100×100$\times$100 × longer than points along its front or back. Specifically, points along the sides are pulled upward by the liquid-vapor interface for a time on the order of $l/v$, from roughly 1 to 100 s. Points at the front and back are only pulled upward for a time on the order of $\gamma/Ev$, from roughly 0.01 to 1 s. The relaxation data in Fig. 4(g) supports this hypothesis, as rails formed at lower speeds (i.e., longer contact times) are taller.
The growth and relaxation of the elastocapillary rails are dependent on the mechanism of dissipation in the substrate. This may be viscoelastic or poroelastic in origin. Previous work [38, 21] has shown that poroelasticity can lead to ridge decay times comparable to the contact time between the droplet and substrate. While the viscoelastic timescale ($\tau\sim 10^{-1}$ s) governs the short-time relaxation of a wetting ridge, the poroelastic timescale ($\tau_{\mathrm{p}}$) governs the long-time relaxation, scaling as $w\gamma/(ED^{*})\sim 1$ s, where $D^{*}\sim 10^{-8}~{}\mathrm{m^{2}/s}$ is the poroelastic diffusion coefficient [21]. Interestingly, the time for which points along the sides are pulled upward by the liquid-vapor interface (roughly 1 to 100 s) is comparable to this poroelastic timescale, while the time for which points at the front and back are pulled upward (roughly 0.01 to 1 s) is closer to the viscoelastic relaxation time. Consequently, due to poroelasticity, elastocapillary ridges formed along the direction of motion would be more persistent than those at the leading and trailing edges of the droplet, enhancing the formation of sheddable elastocapillary rails.
@@
The poroelastic argument suggests a potential effect of droplet size on the rails: for larger droplets, the longer pull-up time could lead to slower rail decay [21]. Investigating this size dependence is an interesting topic for future research. Additionally, the influence of poroelasticity on dissipation remains an intriguing open question to explore in future studies.